Check whether following system of equations has a solution.
3x+2y+6z+4w=4
x+2y+2z+w=5
x+z+3w=3
1
Expert's answer
2020-03-29T10:00:32-0400

First we construct and transform a matrix for this system:

"\begin{pmatrix}
 3 & 2 & 6 & 4 &|& 4 \\
 1 & 2 & 2 & 1 &|& 5 \\
 1 & 0 & 1 & 3 &|& 3 
\end{pmatrix} \overset{R_1 - 3R_2, R_2 - R_3}{=}
\begin{pmatrix}
 0 & -4 & 0 & 1 &|& -11 \\
 0 & 2 & 1 & -2 &|& 2 \\
 1 & 0 & 1 & 3 &|& 3 
\end{pmatrix} \overset{R_3 \to R_1, R_1 \to R_3}{=} \\
= \begin{pmatrix}
 1 & 0 & 1 & 3 &|& 3 \\
 0 & 2 & 1 & -2 &|& 2 \\
 0 & -4 & 0 & 1 &|& -11 
\end{pmatrix} \overset{R_3 + 2R_2}{=}
\begin{pmatrix}
 1 & 0 & 1 & 3 &|& 3 \\
 0 & 2 & 1 & -2 &|& 2 \\
 0 & 0 & 2 & -3 &|& -7 
\end{pmatrix}\\"

"z = \frac{3w - 7}{2}\\
y = \frac{-z + 2w + 2}{2} = -\frac{z}{2} + w + 1 = \frac{7 - 3w}{4} + w + 1 = \frac{w + 11}{4}\\
x = -z - 3w + 3 = \frac{7 - 3w}{2} -3w + 3 = \frac{13 - 9w}{2}"

So a solution of the system exists.